OpenStudy (anonymous):

What is the 9th term of the geometric sequence 4, -20, 100, ...?

hartnn (hartnn):

first term =a1 = 4 common ratio = r = -20/4 = -5 use the formula : \(a_n =a_1 r^{n-1} \) where , r =-5, n= 9